### v1

@@ 3.4 Parameter Updates
To investigate language-specific encoding in LLMs, we analyze parameter updates during fine-tuning and compare them across languages to determine whether certain components of the model specialize in processing linguistic information. Meng et al. (2022) suggest that MLP modules primarily store knowledge, while attention modules control information retrieval and selection. SFT models correctly answer approximately 40% of evaluation questions in all languages. However, they require fine-tuning to improve their ability to select and output the correct information. As a result, we expect substantial modifications in the attention modules, particularly in the final layers, while changes in the MLP modules remain limited. Since these datasets differ only in language, not in task or knowledge, analyzing the model updates allows us to pinpoint which layers or components are most crucial for learning language-specific representations.
To compute parameter update, we follow Bandarkar et al. (2024) and calculate the average parameter modifications for each module in each layer. Denoting the pre-trained weight matrix as $W_{p}$ and the fine-tuned weight matrix as $W_{f}$, the average magnitude of differences is given by:
$\Delta W=\frac{1}{n}\sum_{i=1}^{n}|W_{p}^{(i)}-W_{f}^{(i)}|\vspace{-2mm}$ (2)
Figure 3: Parameter updates of monolingual finetuning for Llama-1B, Llama-3B, and Llama-8B.
The results for four languages are shown in Figure 3, with the remaining three languages in Figure 6. As expected, significant modifications occur in the attention modules of the final six layers for Llama-3.2-1B and the final 14 layers for Llama-3.2-3B and Llama-3.1-8B models across all languages. However, in Llama-3.2 models, we observe substantial changes in the MLP modules in these layers for all languages except English, suggesting that these variations are tied to language-specific processing rather than task-related learning. Surprisingly, for Llama-3.1-8B, even the model fine-tuned on English shows a high rate of change similar to other languages. Considering the unexpectedly low accuracy of the Llama-3.1-8B pre-trained model across all languages compared to Llama-3.2-3B, this larger modification could be related to learning the task or acquiring new knowledge rather than just language adaptation.
@@ 4.1 Correct Concept in Correct Language
According to Dubey et al. (2024), only 8% of the pre-training data used for Llama-3 models is multilingual, while the rest is dominated by English general knowledge, mathematics, and code. This suggests a strong bias toward English. Given this imbalance, we hypothesize that the observed MLP module changes in non-English languages may not indicate new knowledge acquisition but rather adjustment in language selection during response generation. Supporting this, Chirkova and Nikoulina (2024) found that when Llama-2-13B is instruction-tuned in English and tested in other languages, it generates responses in a different language from input language in over 30% of cases, with this behavior influenced by training hyperparameters.
@@
$\begin{split}\text{CoCo-CoLa}(L_{i})&\\ &=\frac{|C_{L_{i}\rightarrow L_{i}}-\bigcup\limits_{L_{o}\neq L_{i}}C_{L_{i}\rightarrow L_{o}}|}{|C_{L_{i}\rightarrow L_{i}}~{}\Delta\bigcup\limits_{L_{o}\neq L_{i}}C_{L_{i}\rightarrow L_{o}}|}\end{split}$ (3)
Language Metric 1B 3B 8B
$PLM$ $\rightarrow en$ $\rightarrow L_{i}$ $PLM$ $\rightarrow en$ $\rightarrow L_{i}$ $PLM$ $\rightarrow en$ $\rightarrow L_{i}$
French Cum. Acc. 12.07 52.66 55.73 20.57 62.55 52.97 12.89 58.64 66.01
CoCo-CoLa 49.42 13.47 88.58 52.51 14.73 89.45 58.11 12.32 87.54
@@
CoCo-CoLa 53.87 10.50 91.02 56.53 19.64 89.26 57.49 11.03 87.21
@@
CoCo-CoLa 52.28 12.01 91.24 61.68 15.84 89.18 61.98 9.40 91.35
Table 2: CoCo-CoLa ratio and cumulative accuracy of pretrained model (PLM), English-trained model ($\rightarrow en$), and $L_{i}$-trained model ($\rightarrow L_{i}$) across languages for Llama-1B, Llama-3B, and Llama-8B
The denominator uses the symmetric difference between $C_{L_{i}\rightarrow L_{i}}$ and correct answers in other languages because many answers involve named entities, such as well-known places, books, and individuals. Since most of the languages, use similar scripts, named entities often appear in identical forms across multiple languages. This redundancy leads to overlap between $C_{L_{i}\rightarrow L_{i}}$ and $\bigcup_{L_{o}\neq L_{i}}C_{L_{i}\rightarrow L_{o}}$, which the symmetric difference helps mitigate by ensuring that shared named entities do not artificially inflate the metric.
@@
To evaluate language adherence and accuracy, we pass the input in $L_{i}$ to pre-trained, en-tuned, and $L_{i}$-tuned models. We then compute the CoCo-CoLa ratio and the cumulative accuracy, defined as the proportion of correct answers either in $L_{i}$ or English. The results, presented in Table 2, show that while the en-tuned models and the $L_{i}$-tuned models achieve similar cumulative accuracy on $L_{i}$ input, the CoCo-CoLa ratio is significantly lower for the en-tuned model. This suggests that although the en-tuned model can correctly process the question in $L_{i}$ and retrieve the correct answer at the same rate as the $L_{i}$-tuned model, it frequently generates the answer in English instead of $L_{i}$. Furthermore, analyzing the CoCo-CoLa ratio of the pre-trained model reveals that the model already exhibits a bias toward generating English responses, though this bias is less pronounced than in the en-tuned model. These findings support our hypothesis that the varying rate of parameter updates across languages is related to output language preference. Since the model is already inherently biased toward English, en-tuned results in the least change.
@@
4.2 Partial Training for Efficient Adaptation
In this section, we aim to disentangle task learning from output generation in language $L_{i}$. Our previous results reveal two key observations. First, as shown in Section 4.1, both the en-tuned model and the $L_{i}$-tuned model achieve the same cumulative accuracy on $L_{i}$, indicating that they learn the task equally well. The only difference is their CoCo-CoLa score, meaning that while both models understand the task to the same degree, they generate outputs in different languages. Second, from Section 3.4, we observed that the en-tuned and $L_{i}$-tuned models undergo different parameter updates. Some of these updates are necessary for learning the task itself, while others are specifically related to generating responses in the correct language.
Based on these observations, we hypothesize that fine-tuning specific layers of an en-tuned model on $L_{i}$ can enable it to generate responses in $L_{i}$ without requiring full model updates. Specifically, these layers correspond to the parameters that were updated in the $L_{i}$-tuned model but not in the en-tuned model. To test this hypothesis, we first identify the layers that undergo language-specific updates. We then fine-tune only these layers in the en-tuned model and compare the results to fine-tuning non-specific layers. This comparison allows us to isolate the parameters responsible for output language.
@@ 4.2 Partial Training for Efficient Adaptation
We select layers for partial training based on the variation in parameter update rates observed in Section 3.4. For the Llama-3.2-1B model, we train three variants by unfreezing different sets of layers: (1) layers 11-16, (2) layers 1-5 (chosen to match the parameter count of the final six layers), and layers 1-10 (including all parameters except the final six). We expect the first variant, which targets the final six layers, to be the most language-related and to result in the largest improvement in the CoCo-CoLa ratio, while the other two should have a smaller effect. Similarly, for Llama-3.2-3B, we train two variants by unfreezing layers 15–27 and layers 1–14, again expecting the first variant to be more strongly related language generation. For Llama-3.1-8B, which does not show clear variations in update rates across languages (as noted in Section 3.4), we instead select layers based on the most updated MLP modules. Specifically, we choose layers 16–31 and layers 1–15 for partial training to determine which part of the model is more responsible for language generation. Through this analysis, we aim to verify whether the final layers play a greater role in controlling the output language
@@
As shown in Figure 4, among the partially trained models, training the final layers yields the highest accuracy and CoCo-CoLa ratio for Llama-3.2 models, indicating that these layers play a crucial role in determining output language. Moreover, the accuracy of this partially trained model approaches that of the fully $L_{i}$-tuned model, suggesting that the earlier layers already provide sufficient information for answering questions, even without exposure to $L_{i}$ during training. Interestingly, Hindi, which initially exhibited lower performance than other languages, benefits from cross-lingual transfer, achieving better results with partial training than full training in both Llama-3.2 models. Llama-3.2-3B shows even stronger cross-lingual transfer, improving accuracy for Italian and Portuguese as well. For Llama-3.1-8B, training the second half of the model leads to the best CoCo-CoLa ratio, but the difference in $L_{i}$ accuracy across partial training configurations is less pronounced than in Llama-3.2 models. This model also shows improved accuracy in partial training over full training for German, Italian, and Portuguese.
@@
These findings confirm the hypothesis that the final layers are strongly linked to output language selection. Additionally, for low-resource languages, partially training only the final layers of an en-tuned model can achieve similar or even better accuracy compared to full fine-tuning in the target language. Beyond its effectiveness, partial training is significantly more efficient, reducing training time to half and memory usage to 65% of full training. Furthermore, the model achieves higher accuracy in fewer training steps, requiring less than one epoch, meaning it is trained on fewer data points.
@@ 5 Conclusion
To address this problem, we leveraged insights from parameter updates and CoCo-CoLa results to develop a partial training method that improves language adherence in English-trained models. Our analysis demonstrated a more efficient alternative to full fine-tuning, achieving comparable or even superior performance while significantly reducing the number of updated parameters. In addition, given the widespread availability of instruction-tuned and task-specific English models, partial training of final layers presents a fast and efficient approach to adapting LLMs to new languages.
Limitations
We acknowledge that training hyperparameters can significantly influence the linguistic bias of fine-tuned models, as highlighted by Chirkova and Nikoulina (2024). For instance, while smaller learning rates may reduce bias, they can also lead to degraded task performance. Due to resource constraints, we used a single set of hyperparameters optimized for task performance rather than systematically exploring bias mitigation strategies. Additionally, we applied the same hyperparameter settings across all languages and model sizes, though fine-tuning them individually for each model-language pair could potentially yield better results.
@@
Another key limitation is that our study focuses on languages from the same language family, which are relatively close to each other and exhibit significant token overlap, facilitating cross-lingual transfer. The models we evaluated were also trained on a limited set of languages with similar characteristics. None of the studied languages fall into the low-resource category, meaning our findings may not generalize to massively multilingual models trained on a more diverse set of languages. Finally, our method is restricted to languages present in the model’s pretraining data and cannot be applied to completely new languages.
Ethical Statement
This research investigates language adherence in multilingual large language models (LLMs) and proposes partial training methods for efficient adaptation. Our work aims to enhance linguistic fairness and accessibility by mitigating biases that favor high-resource languages. We acknowledge that training data composition and fine-tuning decisions can introduce unintended biases, which may disproportionately affect underrepresented languages. While our findings contribute to more equitable multilingual model adaptation, they are limited to languages present in the model’s pretraining data and may not generalize to unseen languages.
Additionally, we use machine translation tools (e.g., Google Translate) to ensure consistency in question-answer pairs. While back-translation was employed to verify accuracy, potential translation artifacts could impact evaluation. We encourage further work to assess our method’s applicability to a broader set of languages, particularly low-resource and non-Indo-European languages.
@@
Appendix A Appendix
A.1 Implementation details

### v2

@@ 3.4 Parameter Updates
To investigate language-specific encoding in LLMs, we analyze parameter updates during fine-tuning and compare them across languages to determine whether certain components of the model specialize in processing linguistic information. Meng et al. [17] suggest that MLP modules primarily store knowledge, while attention modules control information retrieval and selection. SFT models correctly answer approximately 40% of evaluation questions in all languages. However, they require fine-tuning to improve their ability to select and output the correct information. As a result, we expect substantial modifications in the attention modules, particularly in the final layers, while changes in the MLP modules remain limited. Since these datasets differ only in language, not in task or knowledge, analyzing the model updates allows us to pinpoint which layers or components are most crucial for learning language-specific representations.
To compute parameter update, we follow Bandarkar et al. [32] and calculate the average parameter modifications for each module in each layer. Denoting the pre-trained weight matrix as $W_{p}$ and the fine-tuned weight matrix as $W_{f}$, the average magnitude of differences is given by:
$\Delta W=\frac{1}{n}\sum_{i=1}^{n}|W_{p}^{(i)}-W_{f}^{(i)}|$ (2)
Figure 3: Heatmaps of parameter update magnitudes during monolingual fine-tuning on English (top) and French (bottom) across different LLMs.
The results for English and French are shown in Figure 3, with the remaining languages in Figure 6. As expected, significant modifications occur in the attention modules of the final six layers for Llama-1B and the final 14 layers for Llama-3B, Llama-8B, and Gemma-4B models across all languages. However, in Llama-3.2 models and Gemma-4B model, we observe substantial changes in the MLP modules in these layers for all languages except English, suggesting that these variations might be tied to language-specific processing rather than task-related learning. Surprisingly, for Llama-8B, even the model fine-tuned on English shows a high rate of change similar to other languages. Considering the unexpectedly low accuracy of the Llama-8B pre-trained model across all languages compared to Llama-3B, this larger modification could be related to learning the task or acquiring new knowledge rather than just language adaptation.
@@ 4.1 Correct Concept in Correct Language
According to Dubey et al. [40], only 8% of the pre-training data used for Llama-3 models is multilingual, while the rest is dominated by English general knowledge, mathematics, and code. This suggests a strong bias toward English. Given this imbalance, we hypothesize that the observed MLP module changes in non-English languages may not indicate new knowledge acquisition but rather adjustment in language selection during response generation. Supporting this, Chirkova and Nikoulina [2] found that when Llama-2-13B is instruction-tuned on English and tested in other languages, it generates responses in a different language from input language in over 30% of cases, with this behavior influenced by training hyperparameters.
@@
Table 2: CoCo-CoLa ratio (Ratio) and cumulative accuracy (Acc) of pretrained model (PLM), English-tuned model ($\rightarrow en$), and $L_{i}$-tuned model ($\rightarrow L_{i}$) across languages for Llama-1B, Llama-3B, Llama-8B, and Gemma-4B.
Language Metric Llama-1B Llama-3B Llama-8B Gemma-4B
$PLM$ $\rightarrow en$ $\rightarrow L_{i}$ $PLM$ $\rightarrow en$ $\rightarrow L_{i}$ $PLM$ $\rightarrow en$ $\rightarrow L_{i}$ $PLM$ $\rightarrow en$ $\rightarrow L_{i}$
French Acc 12.07 52.66 55.73 20.57 62.55 52.97 12.89 58.64 66.01 18.67 65.16 63.23
Ratio 49.42 13.47 88.58 52.51 14.73 89.45 58.11 12.32 87.54 50.77 19.22 90.14
@@
Ratio 56.40 12.73 91.07 63.37 15.99 85.10 51.41 11.49 90.73 51.89 14.98 90.69
[…]
Ratio 52.28 12.01 91.24 61.68 15.84 89.18 61.98 9.40 91.35 51.15 14.70 91.36
The denominator uses the symmetric difference between $C_{L_{i}\rightarrow L_{i}}$ and correct answers in other languages because many answers involve named entities, such as well-known places, books, and individuals. Since most of the languages in this work use similar scripts, named entities often appear in identical forms across multiple languages. This redundancy leads to overlap between $C_{L_{i}\rightarrow L_{i}}$ and $\bigcup_{L_{o}\neq L_{i}}C_{L_{i}\rightarrow L_{o}}$, which the symmetric difference helps mitigate by ensuring that shared named entities do not artificially inflate the metric.
@@
To evaluate language adherence and accuracy, we pass the input in $L_{i}$ to pre-trained, en-tuned, and $L_{i}$-tuned models. We then compute the CoCo-CoLa ratio and the cumulative accuracy, defined as the proportion of correct answers either in $L_{i}$ or English. The results, presented in Table 2, show that while the en-tuned models and the $L_{i}$-tuned models achieve comparable cumulative accuracy on $L_{i}$ input, the CoCo-CoLa ratio is significantly lower for the en-tuned model. This suggests that although the en-tuned model can correctly process the question in $L_{i}$ and retrieve the correct answer at the same rate as the $L_{i}$-tuned model, it frequently generates the answer in English instead of $L_{i}$. Furthermore, analyzing the CoCo-CoLa ratio of the pre-trained model reveals that the model already exhibits a bias toward generating English responses, though this bias is less pronounced than in the en-tuned model. These findings support our hypothesis that the varying rate of parameter updates across languages is related to output language preference. Since the model is already inherently biased toward English, en-tuned results in the least MLP change compared to other languages.
4.2 Partial Training for Language Adaptation
In this section, we aim to disentangle task learning from output generation in language $L_{i}$. Our previous results reveal two key observations. First, as shown in Section 4.1, both the en-tuned model and the $L_{i}$-tuned model achieve comparable cumulative accuracy on $L_{i}$, indicating that they learn the task equally well. The only difference is their CoCo-CoLa score, meaning that while both models understand the task to the same degree, they generate outputs in different languages. Second, from Section 3.4, we observed that the en-tuned and $L_{i}$-tuned models undergo different parameter updates. Some of these updates are necessary for learning the task itself, while others may specifically steer the model toward producing responses in the intended language.
@@ 4.2 Partial Training for Language Adaptation
Based on these observations, we hypothesize that fine-tuning specific layers of an en-tuned model on $L_{i}$ can enable it to generate responses in $L_{i}$ without requiring full model updates. Specifically, these layers correspond to the parameters that were updated in the $L_{i}$-tuned model but not in the en-tuned model. To test this hypothesis, we first identify the layers that undergo language-specific updates. We then fine-tune only these layers in the en-tuned model and compare the results to fine-tuning other layers. This comparison allows us to isolate the parameters responsible for output language.
@@
We select layers for partial training based on the variation in parameter update rates observed in Section 3.4. For the Llama-1B model, we train three variants by unfreezing different sets of layers: (1) layers 11-16, (2) layers 1-5 (chosen to match the parameter count of the final six layers), and layers 1-10 (including all parameters except the final six). We expect the first variant to be the most language-related and to result in the largest improvement in the CoCo-CoLa ratio, while the other two should have a smaller effect. For Llama-3B and Gemma-4B, we similarly train two variants each: unfreezing layers 15–28 and 1–14 for Llama-3B, and layers 21–34 and 1–20 for Gemma-4B. Again, we expect the final-layer variants to have a stronger relationship to language generation. For Llama-8B, which does not show clear variations in update rates across languages (as noted in Section 3.4), we instead select layers based on the most updated MLP modules. Specifically, we choose layers 16–32 and layers 1–15 for partial training to determine which part of the model is more responsible for language generation. Through this analysis, we aim to verify whether the final layers play a greater role in controlling the output language
@@
As shown in Figure 4, among the partially trained models, unfreezing the final layers results in the highest accuracy and CoCo-CoLa ratio for all models, highlighting the crucial role these layers play in determining the output language. Notably, the accuracy of this partially trained configuration closely approaches that of the fully $L_{i}$-tuned model, suggesting that the earlier layers already encode sufficient information for question answering, even without direct exposure to $L_{i}$ during training. Interestingly, Hindi-which initially exhibited lower performance than other languages-benefits significantly from cross-lingual transfer, achieving better results with partial training than with full training in both Llama-3.2 models. Llama-3B demonstrates even stronger cross-lingual transfer, with improved accuracy for Italian and Portuguese as well. For Llama-8B and Gemma-4B, training the second half of the model yields the highest CoCo-CoLa ratio; however, the differences in $L_{i}$ accuracy across partial training configurations are less pronounced than in the Llama-3.2 models. These models also show improved accuracy with partial training compared to full training for German, Italian, and Portuguese in Llama-8B, and for French, Portuguese, and Spanish in Gemma-4B.
@@
For low-resource languages, partially training only the final layers of an en-tuned model can achieve similar or even better accuracy compared to full fine-tuning in the target language. Beyond its effectiveness, partial training is significantly more efficient, reducing training time to half and memory usage to 65% of full training. Furthermore, the model achieves higher accuracy in fewer training steps, requiring less than one epoch, meaning it is trained on fewer data points.
These findings confirm the hypothesis that the final layers are linked to output language selection, whereas initial and middle layers have less effect on the output language. Our results are aligned with concurrent work suggesting that LLMs process input in three stages: understanding the input, reasoning and knowledge retrieval in a shared space among languages, and generating output [45, 46, 47]. Although it remains debated whether this shared knowledge space is language agnostic [46] or whether the model simply thinks in English [45, 47], these works, alongside ours, all suggest that the process happening in middle layers is not dependent on the input language. However, what previous works overlooks is that the final stage is defective and cannot generate the response in the correct language. We believe this phenomenon has led to misleading evaluations and the belief that multilingual LLMs think better in English [48]. Our work emphasizes the importance of considering both correctness and language adherence, as relying on output accuracy against the ground truth does not provide a complete picture of a model’s ability to reason and operate in non-dominant languages.
4.3 Improving Language Adherence in Multilingual LLMs
Table 3: CoCo-CoLa ratio (Ratio) and cumulative accuracy (Acc) of models partially trained on balanced multilingual data, with averages across all languages.
French German Hindi Italian Portuguese Spanish Average
Model Ratio Acc Ratio Acc Ratio Acc Ratio Acc Ratio Acc Ratio Acc Ratio Acc
Llama-1B 82.15 47.53 64.47 39.02 90.75 28.00 83.58 50.39 73.89 41.51 72.72 41.65 77.93 41.35
Llama-3B 78.37 42.19 79.55 36.14 85.34 33.54 80.11 41.61 74.27 49.04 78.92 44.22 79.43 41.12
Llama-8B 75.95 67.62 85.29 49.75 96.89 32.00 87.83 59.77 88.63 64.55 86.94 38.01 86.92 51.95
Gemma-4B 88.35 57.38 88.04 55.88 83.71 37.32 90.22 64.21 87.00 60.69 87.34 61.81 87.45 56.22
As demonstrated in Section 4.1, multilingual LLMs exhibit a strong linguistic bias toward English, the most prevalent language in their training data. In Section 4.2, we further established that this bias is closely linked to the model’s final layers. To investigate whether this bias can be mitigated and to enable the model to better adhere to the input language, we take the en-tuned model and, rather than adapting it to a single target language, we partially fine-tune the language-related layers using a balanced multilingual dataset, where all languages appear with equal frequency in the training data.
As shown in Table 3, the average CoCo-CoLa ratio for multilingually fine-tuned Gemma-4B and Llama-8B reaches 87.45% and 86.92%, respectively, while Llama-1B and Llama-3B achieve slightly lower ratios of 77.93% and 79.43%. These results are similar to the monolingual models partially trained for each language (Section A.6). These findings indicate that, even when starting from a model pretrained on biased data, fine-tuning only the final layers on a balanced multilingual dataset substantially improves language adherence across all languages. Notably, for Llama-8B and Gemma-4B, the accuracy of the resulting multilingual model is competitive with models fully fine-tuned for each individual language, despite using only 200 datapoints per language during training.
@@ 5 Conclusion
To address this problem, we leveraged insights from parameter updates and CoCo-CoLa results to develop a partial training method that improves language adherence in English-trained models. Our analysis demonstrated a more efficient alternative to full fine-tuning, achieving comparable or even superior performance while significantly reducing the number of updated parameters. Additionally, we showed that partial training on balanced multilingual data achieves similar language adherence to monolingual training. Given the widespread availability of instruction-tuned and task-specific English models, partial training of final layers presents a fast and efficient approach for improving language adherence and adapting LLMs to new languages.
Appendix A Appendix
A.1 Limitations
We acknowledge that training hyperparameters can influence the linguistic bias of fine-tuned models, as highlighted by Chirkova and Nikoulina [2]. For instance, while smaller learning rates may reduce bias, they can also lead to degraded task performance. Due to resource constraints, we used a single set of hyperparameters optimized for task performance. Additionally, we applied the same hyperparameter settings across all languages and model sizes, though fine-tuning them individually for each model-language pair could potentially yield better results.
@@ A.1 Limitations
Another limitation is that our study focuses on languages that mostly come from the same language family, and are relatively close to each other. As a result these languages exhibit significant token overlap, facilitating cross-lingual transfer. The models we evaluated were also trained on a limited set of languages with similar characteristics. The studied languages mainly fall into the mid- or high-resource category, meaning our findings may not generalize to massively multilingual models trained on a more diverse set of languages.
A.2 Ethical Statement
This research investigates language adherence in multilingual large language models and proposes partial training methods for efficient adaptation. Our work aims to enhance linguistic fairness and accessibility by mitigating biases that favor high-resource languages. We acknowledge that training data composition and fine-tuning decisions can introduce unintended biases, which may disproportionately affect underrepresented languages. While our findings contribute to more equitable multilingual model adaptation, they are limited to languages present in the model’s pretraining data and may not generalize to unseen languages. We encourage further work to assess our method’s applicability to a broader set of languages, particularly low-resource and non-Indo-European languages.
@@ A.2 Ethical Statement
A.3 Implementation details